OpenStudy (anonymous):

in a given day the mill turns out three times as many units of plywood as lumber. it makes a profit of $30 on a unit of lumber and $40 on a unit of plywood. how many of each unit must be produced and sold in order to make a profit of $15300

OpenStudy (anonymous):

Let L = the number of units of lumber and let p = the number of plywood units. Then P = 3L because you would have to make 3 times the number of lumber for them to be exactly the same in number. Now you need a cost equation. Customers need to buy a co,300mbination of $30L plus $40P added together to total $15,300. It would have to look like: 30L + 40P = 15,300 Substitute P = 3L into the last equation and solve for L. One you have L and you substitute that in the P = 3L equation to find P.

OpenStudy (anonymous):

i still cant get it

OpenStudy (anonymous):

Define a production unit as 1 lumber unit and 3 plywood units. Divide the profit from one production unit, $30 + 3 * $40, $150, into $15300 the expected profit. The result is 102 production units are required. From the composition of a production unit it follows that 102 and 306, lumber and plywood units, have to be manufactured respectively.

OpenStudy (anonymous):

Plug the P = 3L into where the P is in the equation 30L + 40P = 15,300 that is: 30L + 40(3L) = 15, 300 Now solve for L.